OpenStudy (anonymous):

What is the value of the y variable in the solution to the following system of equations? 5x + 3y = 7 3x - 5y = -23 -1 1 -4 4

OpenStudy (jazzyfa30):

give me one min

OpenStudy (anonymous):

alright

OpenStudy (anonymous):

5x + 3y = 7 Solving for variable 'x'. Move all terms containing x to the left, all other terms to the right. Add '-3y' to each side of the equation. 5x + 3y + -3y = 7 + -3y Combine like terms: 3y + -3y = 0 5x + 0 = 7 + -3y 5x = 7 + -3y Divide each side by '5'. x = 1.4 + -0.6y Simplifying x = 1.4 + -0.6y now for 3x - 5y = -23 3x + -5y = -23 Solving for variable 'x'. Move all terms containing x to the left, all other terms to the right. Add '5y' to each side of the equation. 3x + -5y + 5y = -23 + 5y Combine like terms: -5y + 5y = 0 3x + 0 = -23 + 5y 3x = -23 + 5y Divide each side by '3'. x = -7.666666667 + 1.666666667y Simplifying x = -7.666666667 + 1.666666667y

OpenStudy (anonymous):

Seriously?

OpenStudy (jazzyfa30):

even im confusd

OpenStudy (anonymous):

Yeah.. It has to be one of the answers above

OpenStudy (anonymous):

lol well i sovled both am i suppose to do this? or that 3*5x + 3y*3 = 7*3 -3*3x + -5y*-3 = -23*-3 ????? den mover or to ur right?? sum like dat?

OpenStudy (jazzyfa30):

so lets start with the first equation do you know how to turn it into slope intercept form y=mx+b

OpenStudy (anonymous):

yes. add 5x to both sides?

OpenStudy (anonymous):

or 3y to both sides

OpenStudy (anonymous):

yes 3y to both sides

OpenStudy (anonymous):

Okay so now it would be 5x = 7 + 3y

OpenStudy (anonymous):

ur solving for y right so other way around if ur tyrna elimnate y but here Add '-5x' to each side of the equation. 5x + -5x + 3y = 7 + -5x Combine like terms: 5x + -5x = 0 0 + 3y = 7 + -5x 3y = 7 + -5x

OpenStudy (jazzyfa30):

you make things complicated @dopekidsimmy

OpenStudy (anonymous):

I dont think it becomes 7 + -5x

OpenStudy (anonymous):

Maybe 7 - 5x?

OpenStudy (jazzyfa30):

it would be 3y=5x+7

OpenStudy (anonymous):

and what happenes when u add a positive ad a negative

OpenStudy (anonymous):

and a "

OpenStudy (jazzyfa30):

my b its 3y=-5x+7

OpenStudy (anonymous):

Okay

OpenStudy (jazzyfa30):

so divide by 3 on both sides what do you gt

OpenStudy (anonymous):

I dont see how you can divide that at this point

OpenStudy (anonymous):

but i seem like its um solve im confused my self!? wowwww

OpenStudy (jazzyfa30):

y=-1.6x+2.3

OpenStudy (anonymous):

Im confused.. None of this adds up to the answers above. I'll just have to ask someone else. But thanks for your help. ps. if you dont know how to do this then you shouldnt be trying to help people on here.

OpenStudy (anonymous):

Idk who to give the medal to

OpenStudy (jazzyfa30):

np

OpenStudy (anonymous):

give it to jazzy she derserves it im confused

OpenStudy (jazzyfa30):

awwww thanks @dopekidsimmy and thanks @SavannahSaturn

OpenStudy (anonymous):

lol no prob @jazzyfa30

OpenStudy (anonymous):

My pleasure. and again I wasnt trying to be rude =)

OpenStudy (anonymous):

oh my bad its open in another screen lol

OpenStudy (anonymous):

butt hey!!!!!!!!! hello!!!!! anyone!!!! hello!!! testing!!! 1 2 3 !!!!

OpenStudy (anonymous):

Oh hey mate :)